# Calc-CMU at SemEval-2024 Task 7: Pre-Calc – Learning to Use the Calculator Improves Numeracy in Language Models

### v2

@@ Abstract
Calc-CMU at SemEval-2024 Task 7: Pre-Calc – Learning to Use the Calculator Improves Numeracy in Language Models
Vishruth Veerendranath  and Vishwa Shah  and Kshitish Ghate Carnegie Mellon University {vveerend, vishwavs, kghate}@andrew.cmu.edu
@@ 1 Introduction
We present Pre-Calc, a pre-finetuning objective of learning to use the calculator, to improve numerical abilities in language models. We propose Pre-Calc objectives for both the encoder-only and encoder-decoder classes of language models, and use a combination of the MAWPS Koncel-Kedziorski et al. (2016), SVAMP Patel et al. (2021), AsDiv-A Miao et al. (2020) datasets to pre-finetune the models.
Our encoder-only objective, used to pre-finetune BERT and RoBERTa models, offers quick and efficient processing suitable for tasks where speed is paramount. The Pre-Calc versions of the models show competent performance on all quantitative downstream tasks from NumEval Chen et al. (2023) and substantial improvements on 4 out of 6 sub-tasks with an improvement greater than 10 points for RedditNLI and AWPNLI specifically.
@@ 2.1 Downstream Tasks
We focus on the QNLI and QQA tasks of NumEval Task 1 Chen et al. (2023) as our downstream tasks.
QNLI is the task of making natural language inferences based on quantitative clues. This dataset is adopted from the EQUATERavichander et al. (2019) and is composed of NewsNLI, RedditNLI, AWPNLI and RTE-Quant. StressTest involves numerical reasoning instances from Naik et al. (2018), used as a synthetic sanity check.
QQA involves the task of multiple-choice question answering involving commonsense as well as quantitative comparisons. The dataset for this is adopted from Task 3 of NumGLUE Mishra et al. (2022) and the Quarel dataset Tafjord et al. (2019), which includes questions from quantitative domains such as physics and economics.
@@ 2.2 Pre-Finetuning Data
We use the MAWPS dataset Koncel-Kedziorski et al. (2016), SVAMP Patel et al. (2021) and AsDiv-A Miao et al. (2020) as the numerical domain datasets for pre-finetuning. These consist of simple arithmetic word problems, along with their numerical solutions. The three datasets create a dataset with 4,225 total examples which are challenging and require understanding the context of numbers, represented either as digits or in words.
We construct this dataset from the Calc-X collection Kadlčík et al. (2023) that has been annotated with equations for each problem, as well as <gadget> annotations in the answer to train a model to use a calculator when the <gadget> token is produced. We use the annotations of the equations particularly in our methodology.
@@ 3.1.1 Data Preprocessing
We preprocess Calc-MAWPS, Calc-SVAMP and Calc-AsDiv-A (from the Calc-X collection) Kadlčík et al. (2023) and add 2 new features required for Pre-Calc. First is the operand tag sequence, which is a sequence of binary tags that is 1 if the original token it corresponds to is an operand and 0 if it isn’t. Secondly we extract the Operation, which is the operation among {+ (add), - (subtract), * (multiply), / (divide)} that is required for the question. We extract the operation either directly from the equation or the reasoning chain in Calc-X and generate the operand tag sequence, by first extracting the operands and then tagging the occurances of the operands in the binary sequence with a 1. As part of this process we also filter out instances where there are more than one distinct operations as part of the equation.
@@ 3.2 Encoder-Decoder
Encoder-Decoder or Decoder based models provide the abilities of long-form unbounded generations. This is advantageous for numerical problems, where multiple intermediate operations might be required for computation or reasoningWei et al. (2022). By reframing our task to output expressions, we distil the task to output the set of operations, leaving the computation part to the tool.
@@ 3.2.1 Data Preprocessing
As mentioned earlier, we use the MAWPS dataset for training our model to output expressions. As each instance in MAWPS consists of a question and a single numerical answer, to obtain closer resemblance to NLI format tasks, we reframe the question-numerical answer instance to a pair of complete sentences using prompting with LLaMa-7B Touvron et al. (2023) as this is a simple text generation task. To obtain contradiction pairs, we perturb the true numerical answer by a small value (ranging from -5 to +5) before passing to the LLaMa-7B model as these will create harder instances for the model to learn from. We additionally also use the Multi-NLIWilliams et al. (2018) dataset to retain and improve textual inferential abilities of the model. We train on this combined data as Seq2Seq generation task. Combining these tasks should allow the model to infer both semantic and computational capabilities.
@@ 4.1 Baselines
We compare the performance of our method against several baseline models on tasks that require numeracy. Following Chen et al. (2023), the baseline methods involve reframing techniques, namely Original, Digit-based, and Scientific Notation methods, and are pre-finetuned on the Comparing Numbers Dataset (CND). Each of these methods are applied to both BERT Devlin et al. (2019) and RoBERTa Liu et al. (2019) to create two versions of each baseline method.
@@ 4.2 Encoder-Only
We use the pretrained BERT and RoBERTa base models and pre-finetune as per Section 3.1.2. We use the 4-class cross-entropy loss for training the Operation Classification head, and a 2-class cross-entropy (equivalent to binary cross entropy) loss for the Operand Identification head. The models are trained with the Adam optimizer for 20 epochs, a batch size of 8, and a learning rate of 5e-4. The checkpoint after this pre-finetuning is named Pre-Calc-BERT111 https://huggingface.co/collections/Calc-CMU/pre-calc-657a5ad5f1ae42fb12364563 or Pre-Calc-RoBERTa1.
We then finetune Pre-Calc-BERT and Pre-Calc-RoBERTa on the downstream tasks of QNLI and QQA using the same hyperparameters used by the CN-BERT baselines Chen et al. (2023) — AdamW optimizer with a learning rate of 5e-5, batch size of 8 for 5 epochs. We use 10-fold cross validation to report our results for the tasks where an explicit test split is not available.
@@ 4.3 Encoder-Decoder
We use Flan-T5 as our base model, For pre-finetuning, we collect a balanced sample consisting about 4200 instances created from MAWPS for math-nli task and 3900 instances extracted from Multi-NLI for text-nli task.1
As this is a sequence generation task, the objective is same as that in CausalLM modeling for next-word prediction. We use AdamW optimizer with a learning rate of 5e-5, batch size of 8 for 5 epochs. We do not perform any fine-tuning on our downstream tasks, and show results for prompt based few-shot evaluations on each task. We call our model FlanT5-Pre-Calc1 and use 2 baselines, Flan-T5 few-shot and Flan-T5-ST only with standard text training1.
@@ 5.2 Encoder-Decoder
We present the results for our Encoder-Decoder based approach in Table 2. We see that for AWPNLI which requires explicit computation, FlanT5(Pre-Calc) achives almost double performance compared to FlanT5-few shot and FlanT5-ST, showing that Flan-T5 originally did not have this capability to evaluate expressions and compare values and this property cannot be instilled only via text finetuning as can be seen from the performance of FlanT5-ST. Further compared to prior works Chen et al. (2023), this achieves state-of-the art results on AWPNLI.
@@ 6.2 Operation wise difficulty for FlanT5
Figure 5: Operation wise error for FlanT5-Pre-Calc
@@ 7 Related Work
Numeracy, or the ability to understand and work with numbers, is a critical aspect that has been relatively underexplored compared to other linguistic competencies in NLP models. Spithourakis and Riedel (2018) emphasized the need for LMs to better understand numbers, setting a precedent for subsequent research.
@@
In addition to these studies focused on numeral representation, other researchers have investigated the potential of pre-finetuning tasks to enhance LM capabilities. Aghajanyan et al. (2021) introduced a massive multi-task representation with pre-finetuning, demonstrating the efficacy of pre-finetuning in improving model performance across a range of tasks.
Geva et al. (2020) proposed GENBERT, which is trained on automatically-generated synthetic data in a multi-task setup. This training significantly improves performance on numerical reasoning tasks such as DROP and math word problems, while maintaining high performance on standard reading comprehension tasks. Wang et al. (2017) presented a deep neural solver, a hybrid model combining the RNN with a similarity-based retrieval to translate math word problems into equation templates.

### v3

# Pre-Calc: Learning to Use the Calculator Improves Numeracy in Language Models

@@ Abstract
Machine Learning, ICML
@@ 1 Introduction
We present Pre-Calc, a pre-finetuning objective of learning to use the calculator, to improve numerical abilities in language models. We propose Pre-Calc objectives for both the encoder-only and encoder-decoder classes of language models, and use a combination of the MAWPS (Koncel-Kedziorski et al., 2016), SVAMP (Patel et al., 2021), AsDiv-A (Miao et al., 2020) datasets to pre-finetune the models.
Our encoder-only objective, used to pre-finetune BERT and RoBERTa models, offers quick and efficient processing suitable for tasks where speed is paramount. The Pre-Calc versions of the models show competent performance on 6 quantitative downstream tasks from Chen et al. (2023) and substantial improvements on 4 sub-tasks with an improvement greater than 10 points for RedditNLI and AWPNLI specifically.
@@ 2.1 Downstream Tasks
We focus on downstream quantitative reasoning tasks, specifically QNLI and QQA (Chen et al., 2023).
QNLI is the task of making natural language inferences based on quantitative clues. This dataset is adopted from the EQUATE(Ravichander et al., 2019) and is composed of NewsNLI, RedditNLI, AWPNLI and RTE-Quant. StressTest involves numerical reasoning instances from Naik et al. (2018), used as a synthetic sanity check.
QQA involves the task of multiple-choice question answering involving commonsense as well as quantitative comparisons. The dataset for this is adopted from Task 3 of NumGLUE (Mishra et al., 2022) and the Quarel dataset (Tafjord et al., 2019), which includes questions from quantitative domains such as physics and economics.
@@ 2.2 Pre-Finetuning Data
We use the MAWPS dataset (Koncel-Kedziorski et al., 2016), SVAMP (Patel et al., 2021) and AsDiv-A (Miao et al., 2020) as the numerical domain datasets for pre-finetuning. These consist of simple arithmetic word problems, along with their numerical solutions. The three datasets create a dataset with 4,225 total examples which are challenging and require understanding the context of numbers, represented either as digits or in words.
We construct this dataset from the Calc-X collection (Kadlčík et al., 2023) that has been annotated with equations for each problem, as well as <gadget> annotations in the answer to train a model to use a calculator when the <gadget> token is produced. We use the annotations of the equations particularly in our methodology.
@@ 3.1.1 Data Preprocessing
We preprocess Calc-MAWPS, Calc-SVAMP and Calc-AsDiv-A (from the Calc-X collection) (Kadlčík et al., 2023) and add 2 new features required for Pre-Calc. First is the operand tag sequence, which is a sequence of binary tags that is 1 if the original token it corresponds to is an operand and 0 if it isn’t. Secondly we extract the Operation, which is the operation among {+ (add), - (subtract), * (multiply), / (divide)} that is required for the question. We extract the operation either directly from the equation or the reasoning chain in Calc-X and generate the operand tag sequence, by first extracting the operands and then tagging the occurances of the operands in the binary sequence with a 1. As part of this process we also filter out instances where there are more than one distinct operations as part of the equation.
@@ 3.2 Encoder-Decoder
Encoder-Decoder or Decoder based models provide the abilities of long-form unbounded generations. This is advantageous for numerical problems, where multiple intermediate operations might be required for computation or reasoning(Wei et al., 2022). By reframing our task to output expressions, we distil the task to output the set of operations, leaving the computation part to the tool.
@@ 3.2.1 Data Preprocessing
As mentioned earlier, we use the MAWPS dataset for training our model to output expressions. As each instance in MAWPS consists of a question and a single numerical answer, to obtain closer resemblance to NLI format tasks, we reframe the question-numerical answer instance to a pair of complete sentences using prompting with LLaMa-7B (Touvron et al., 2023) as this is a simple text generation task. To obtain contradiction pairs, we perturb the true numerical answer by a small value (ranging from -5 to +5) before passing to the LLaMa-7B model as these will create harder instances for the model to learn from. We additionally also use the Multi-NLI(Williams et al., 2018) dataset to retain and improve textual inferential abilities of the model. We train on this combined data as Seq2Seq generation task. Combining these tasks should allow the model to infer both semantic and computational capabilities.
@@ 4.1 Baselines
We compare the performance of our method against several baseline models on tasks that require numeracy. Following Chen et al. (2023), the baseline methods involve reframing techniques, namely Original, Digit-based, and Scientific Notation methods, and are pre-finetuned on the Comparing Numbers Dataset (CND). Each of these methods are applied to both BERT (Devlin et al., 2019) and RoBERTa (Liu et al., 2019) to create two versions of each baseline method.
@@ 4.2 Encoder-Only
We use the pretrained BERT and RoBERTa base models and pre-finetune as per Section 3.1.2. We use the 4-class cross-entropy loss for training the Operation Classification head, and a 2-class cross-entropy (equivalent to binary cross entropy) loss for the Operand Identification head. The models are trained with the Adam optimizer for 20 epochs, a batch size of 8, and a learning rate of 5e-4. The checkpoint after this pre-finetuning is named Pre-Calc-BERT or Pre-Calc-RoBERTa1.
We then finetune Pre-Calc-BERT and Pre-Calc-RoBERTa on the downstream tasks of QNLI and QQA using the same hyperparameters used by the CN-BERT baselines (Chen et al., 2023) — AdamW optimizer with a learning rate of 5e-5, batch size of 8 for 5 epochs. We use 10-fold cross validation to report our results for the tasks where an explicit test split is not available.
@@ 4.3 Encoder-Decoder
We use Flan-T5 as our base model, For pre-finetuning, we collect a balanced sample consisting about 4200 instances created from MAWPS for math-nli task and 3900 instances extracted from Multi-NLI for text-nli task1.
As this is a sequence generation task, the objective is same as that in CausalLM modeling for next-word prediction. We use AdamW optimizer with a learning rate of 5e-5, batch size of 8 for 5 epochs. We do not perform any fine-tuning on our downstream tasks, and show results for prompt based few-shot evaluations on each task. We call our model FlanT5-Pre-Calc 1 and use 2 baselines, Flan-T5 few-shot and Flan-T5-ST only with standard text training111 https://huggingface.co/collections/Calc-CMU/pre-calc-657a5ad5f1ae42fb12364563.
@@ 5.2 Encoder-Decoder
We present the results for our Encoder-Decoder based approach in Table 2. We see that for AWPNLI which requires explicit computation, FlanT5(Pre-Calc) achives almost double performance compared to FlanT5-few shot and FlanT5-ST, showing that Flan-T5 originally did not have this capability to evaluate expressions and compare values and this property cannot be instilled only via text finetuning as can be seen from the performance of FlanT5-ST. Further compared to prior works (Chen et al., 2023), this achieves state-of-the art results on AWPNLI.
@@ 6.1 Dual-Objective in Encoder-Only Pre-Calc
Figure 3: Operation wise error for FlanT5-Pre-Calc
@@ 7 Related Work
Numeracy, or the ability to understand and work with numbers, is a critical aspect that has been relatively underexplored compared to other linguistic competencies in NLP models. Spithourakis & Riedel (2018) emphasized the need for LMs to better understand numbers, setting a precedent for subsequent research.
@@
In addition to these studies focused on numeral representation, other researchers have investigated the potential of pre-finetuning tasks to enhance LM capabilities. Aghajanyan et al. (2021) introduced a massive multi-task representation with pre-finetuning, demonstrating the efficacy of pre-finetuning in improving model performance across a range of tasks. Geva et al. (2020) proposed GENBERT, which is trained on automatically-generated synthetic data in a multi-task setup. This training significantly improves performance on numerical reasoning tasks such as DROP and math word problems, while maintaining high performance on standard reading comprehension tasks. Wang et al. (2017) presented a deep neural solver, a hybrid model combining the RNN with a similarity-based retrieval to translate math word problems into equation templates.